OpenStudy (anonymous):

which graph represents the line passing though(4,1) and perpendicular to -2x+y=-3

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

also important information you should memorize parallel lines share the same slope perpendicular lines slopes are negative reciprocals of each other ex. \[\frac{ 4}{3} \rightarrow -\frac{3}{4}\]

OpenStudy (anonymous):

@completeidiot can you help me more bc i could find which one is right

OpenStudy (anonymous):

what is the slope of this equation -2x+y=-3 ?

OpenStudy (anonymous):

2/1... idk honestly

OpenStudy (anonymous):

yes that is the slope

OpenStudy (anonymous):

now what is the slope perpendicular to that slope?

OpenStudy (anonymous):

wait how am i suppost to figure that out

OpenStudy (anonymous):

look at the 2nd post

OpenStudy (anonymous):

ok i did

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its just an example showing you what negative reciprocal means

OpenStudy (anonymous):

ok so what do i need to do to figure the question out

OpenStudy (anonymous):

you need to determine the equation of the perpendicular line

OpenStudy (anonymous):

in order to do that, you must determine the perpendicular slope

OpenStudy (anonymous):

then substitute the slope and the point in to point slope form, convert that into slope intercept form and then determine which graph fits your line equation

OpenStudy (anonymous):

wait so how do you figure out the slope on a graph can you give an example

OpenStudy (anonymous):

if the slope is 3, then the perpendicular slope is -1/3

OpenStudy (anonymous):

can you pls juist give me the answer i just really dont understand this